# OncoReg: Medical Image Registration for Oncological Challenges

### v1

@@ 3.4.3 SynDeeds
The image synthesis approach is framed as an additive restoration problem, defining the clean image as the sum of a degraded image and a variable “noise” component. The loss from network learning is computed as the mean squared error (MSE) between the ground truth residual and the output residual. Therefore, the image synthesis network is trained to learn the residual, instead of reconstructing the clean image directly, as shown below:
@@ 3.4.4 Voxelmorph++
Voxelmorph++ utilises a basic single-scale U-Net (Ronneberger et al., 2015) architecture as its baseline backbone, processing image pairs to extract a shared feature representation. This representation serves as the foundation for predicting displacement fields that aim to align fixed and moving images within a designated region of interest.
A distinctive feature of Voxelmorph++ is its discretised heatmap prediction head, which translates the shared feature vectors into a probabilistic tensor representing the likelihood of displacements at specific keypoints. Leveraging pre-generated automatic correspondences based on Förstner keypoints for self-supervision, the prediction head robustly reduces large deformations and enables discretised integral regression of displacements to directly and explicitly match the keypoint supervision. The heatmap is predicted using a transpose 3D convolution followed by several 3D convolution blocks. The output heatmap tensor is then used to define probabilistic displacements, facilitating a more precise alignment by accounting for local variations and potential discrepancies in displacement estimation.
@@
3.4.5 SSKJLBW: Unsupervised Single-shot Registration
@@
To address the challenges of severe noise interference, large-scale deformations caused by respiratory motion, and limited training data (only 10 cases) in thoracic cone-beam CT (ThoraxCBCT) image registration for clinical radiotherapy, this paper proposes a single-shot unsupervised registration framework. Inheriting the optimization paradigm of GroupRegNet (Zhang et al., 2021), the method achieves robust registration without annotated data through multi-scale feature fusion and dynamic regularization constraints. The core workflow begins with noise-adaptive preprocessing of input images, followed by iterative optimization of deformation fields via a cascaded estimation module, and finally outputs smoothed deformation fields constrained by anatomical masks.
@@
The framework incorporates physics-guided preprocessing to decouple noise from anatomical deformations. Input HU values are clamped to [0,2048] to suppress metal artifacts and extreme noise, while edge regions are zero-padded to mitigate respiratory motion artifacts. A lightweight 3D UNet processes standardized 128×128×64 image pairs, leveraging hierarchical features for deformation estimation—shallow layers enforce smooth large-scale deformations, while deeper layers refine local details.
@@ 4.1 ThoraxCBCT
The quantitative results for the ThoraxCBCT task are summarized in Table 2. While ConvexAdam achieved the highest overall ranking across all evaluation metrics, the official winner of the task was SynDeeds. Since ConvexAdam was provided by the organizers as a baseline method, it was not eligible for ranking. However, the performance of ConvexAdam and SynDeeds was very close across most metrics.
When examining individual performance measures, SynDeeds achieved the highest mean Dice Score, while Fourier-Net obtained the lowest Target Registration Error (TRE). The best Hausdorff Distance 95th percentile (HD95) was observed for SSKJLBW, although this method exhibited a less smooth displacement field.
A more detailed breakdown of Dice coefficients achieved by all methods across different anatomical labels is shown in Fig. 3, while the distribution of Target Registration Errors (TREs) is presented in Fig. 5.
@@ 4.2 OncoReg
For the initial challenge evaluation, we used a standard train-validation-test split. However, since many submitted methods did not rely on training data, we later expanded our evaluation to be more robust and meaningful. After completing additional annotations for all images, we conducted a three-fold cross-validation for all trainable models to ensure a fair and comprehensive comparison.
@@
The quantitative results are summarized in Table LABEL:tab:oncoreg and visualized in Fig.6 and Fig.5.
The challenge was won by the DINO-Reg Ensemble, which outperformed all other methods by achieving the lowest Target Registration Error (TRE) and the highest Dice score. It was closely followed by Voxelmorph++ and ConvexAdam, while DeedsBCV and NiftyReg ranked lowest.
@@
Initial 6.41 11.55 34.38 77.56
DINO-RegEns. 3.94 7.62 63.89 76.59 0.062 $>300$ 0.71
Voxelmorph++ 4.66 8.52 62.36 73.96 0.081 $<60$ 0.66
ConvexAdam 4.16 7.94 60.13 76.54 0.066 $(<5)$ 0.61
SynDeeds 4.51 7.81 59.93 74.85 0.100 x 0.56
DINO-Reg 4.53 8.50 57.64 77.03 0.040 $>300$ 0.54
Fourier-Net 6.01 10.46 60.67 71.91 0.103 $>300$ 0.50
MONAI SegResNet 4.78 8.96 53.63 77.37 0.048 x 0.49
DeedsBCV 7.99 11.78 58.33 71.16 0.154 $<15$ 0.40
Niftyreg 7.98 12.43 37.30 75.15 0.061 $<60$ 0.261
@@ 5.1 Evaluation Metrics and Annotation Challenges
Evaluating image registration remains a complex task, as standard metrics often fail to capture all aspects of registration quality and are highly dependent on annotation accuracy. In this challenge, the primary evaluation metric was the Target Registration Error (TRE), which relies on manually placed anatomical landmarks. However, landmark positioning is inherently uncertain in certain structures; while the carina served as a reliable reference point, the aortic arch and vertebral midpoints were more challenging to pinpoint precisely. To minimize inter-observer variability, all landmarks within a given image pair were annotated by the same medical expert, ensuring internal consistency.
Significant effort was dedicated to refining these annotations, including extending the publicly available 4D-Lung dataset and comprehensively annotating our private dataset. This investment considerably enhanced the dataset’s usability. However, annotation challenges were particularly pronounced in cone-beam CT (CBCT) images, where lower resolution, poor contrast, and strong artifacts complicated precise landmark placement. In these cases, tools like TotalSegmentator proved invaluable, significantly reducing annotation time while maintaining quality in a semi-automated setting.
@@ 5.2 Challenge Design and Implementation
Participants had the option to submit results via an automatic evaluation platform on Grand Challenge for the ThoraxCBCT task, but this feature was rarely utilized. Instead, most teams preferred local evaluation. In contrast, the provided Docker example was widely adopted, particularly for handling data loading and saving, which streamlined debugging in the Type 3 challenge setup. Despite these advantages, some recurring issues were identified that should be addressed in future Type 3 challenges.
Organizing such a challenge demands substantial time and resources, but efficient communication and accessibility between organizers and participants played a crucial role in ensuring a smooth workflow. Future iterations should build on this experience to optimize logistical aspects and further reduce participation barriers.
@@ 5.3 Key Findings and Methodological Insights
A key takeaway from the challenge is that the best-performing registration methods leveraged strong feature extraction strategies. The top-performing methods—ConvexAdam and DINO-Reg—demonstrate two distinct but effective approaches:
  • ConvexAdam combines handcrafted and semantic features within a two-stage optimization process. The extracted semantic features play a crucial role in achieving robust registration across different modalities.
@@
Both methods prioritize feature extraction but differ in execution—ConvexAdam explicitly designs features, whereas DINO-Reg learns them in a self-supervised manner.
Beyond these top approaches, an intriguing finding was the potential for further improving ConvexAdam’s performance by integrating it with VoxelMorph++. Specifically, using ConvexAdam-derived Förstner keypoints as input for VoxelMorph++ led to a refined registration process. This effect was particularly evident in OncoReg, where preprocessed keypoints significantly improved VoxelMorph++’s accuracy. However, in ThoraxCBCT, VoxelMorph++ performed notably worse than other methods. Further analysis revealed that its performance was hampered by suboptimal keypoint registration, which proved particularly detrimental in the weakly supervised setting. A post-challenge re-evaluation using keypoints registered by ConvexAdam showed substantial improvement—although this did not affect the official rankings.
Another notable result was the enhancement of DeedsBCV by Li et al., leading to the superior performance of SynDeeds in both registration tasks. This demonstrates that traditional methods, when effectively adapted, can still compete with or even outperform newer deep-learning-based approaches.
@@ Acknowledgments
Data availability
The data for the ThoraxCBCT task is available on learn2reg.grand-challenge.org/oncoreg/. As described earlier the data of the OncoReg task can not be made publicly available, nevertheless, we are open for requests about evaluation on this data that will be computed in-house by us.
A Additional Result Overviews

### v3

@@ 5.3 SynDeeds
The image synthesis approach is framed as an additive restoration problem, defining the clean image as the sum of a degraded image and a variable “noise” component. The loss from network learning is computed as the mean squared error (MSE) between the ground truth residual and the output residual. Therefore, the image synthesis network is trained to learn the residual, instead of reconstructing the clean image directly, as shown in Fig. 3.
@@ 5.4 Voxelmorph++
Voxelmorph++ utilises a basic single-scale U-Net (Ronneberger et al., 2015) architecture as its baseline backbone, processing image pairs to extract a shared feature representation. This representation serves as the foundation for predicting displacement fields that aim to align fixed and moving images within a designated region of interest. A distinctive feature of Voxelmorph++ is its discretised heatmap prediction head, which translates the shared feature vectors into a probabilistic tensor representing the likelihood of displacements at specific keypoints. Leveraging pre-generated automatic correspondences based on Förstner keypoints for self-supervision, the prediction head robustly reduces large deformations and enables discretised integral regression of displacements to directly and explicitly match the keypoint supervision. The heatmap is predicted using a transpose 3D convolution followed by several 3D convolution blocks. The output heatmap tensor is then used to define probabilistic displacements, facilitating a more precise alignment by accounting for local variations and potential discrepancies in displacement estimation.
@@
5.5 TimH: ConstrICON
ConstrICON, introduced by Greer et al. (Greer et al., 2023), is a framework that guarantees inverse consistency by construction. In contrast to other approaches, that utilize penalty terms during training to approximate inverse consistency, ConstrICON achives this property intrinsically through its architectural design and by parameterizing the output transformations of the neural networks by a Lie group. Parameterizing the output of a neural network by a Lie group is assumed to be sufficient, since many types of transformations that are used in medical image registration, such as affine or diffeomorphic transforms, are Lie groups.
Given a neural network $N_{\theta}$ of arbitrary design that takes two input images and whose output can be considered a member of a Lie algebra, a registration network $\Phi$ can be defined, that registers an image $I_{A}$ to an image $I_{B}$:
$\Phi[I_{A},I_{B}]=\exp(N_{\theta}[I_{A},I_{B}]-N_{\theta}[I_{B},I_{A}]).$
The exponential map is implemented using the scaling and squaring method (Arsigny et al., 2006). The registration network $\Phi$ is inverse consistent by construction, as $\Phi[I_{A},I_{B}]\circ\Phi[I_{B},I_{A}]=Id$, where $Id$ denotes the identity transformation. However, the composition of two or more of these networks is not necessarily inverse consistent. To address this issue, Greer et al. proposed the TwoStepConsistent (TSC) operator (Greer et al., 2023), which ensures inverse consistency when combining two networks.
Let $\sqrt{\Phi[I_{A},I_{B}]}$ be the square root of the output transformation of a registration network, defined as $\sqrt{\Phi[I_{A},I_{B}]}\circ\sqrt{\Phi[I_{A},I_{B}]}=\Phi[I_{A},I_{B}]$, where $\Phi$ is trained fulfill $I_{A}\circ\Phi[I_{A},I_{B}]\sim I_{B}$ and let $\Psi$ be a second registration network, that is trained to register $\hat{I}_{A}:=I_{A}\circ\sqrt{\Phi[I_{A},I_{B}]}$ to $\hat{I}_{B}:=I_{B}\circ\sqrt{\Phi[I_{B},I_{A}]}$. Then, the TSC operator is defined as
$TwoStepConsistent\{\Phi,\Psi\}[I_{A},I_{B}]=$
$\sqrt{\Phi[I_{A},I_{B}]}\circ\Psi[\hat{I}_{A},\hat{I}_{B}]\circ\sqrt{\Phi[I_{A},I_{B}]},$
which is inverse consistent if $\Phi$ and $\Psi$ are inverse consistent. Since the TSC operator preserves inverse consistency, it can be recursively applied, thereby extending the procedure to an N-step registration approach that remains inverse consistent.
For the ThoraxCBCT task, we employed three residual U-Nets (Zhang et al., 2018) and applied a four-step preprocessing pipeline. First, images and their corresponding masks were affinely aligned. Then, empty slices were removed, followed by histogram matching. Finally, to reduce GPU memory usage and computational load, images were downscaled to $128\times 128\times 128$ voxels, approximately $47.1\%$ of the original preprocessed size. The square roots of the output transformations of the registration networks were learned implicitly, i.e., we trained an inverse consistent registration network $\tilde{\Phi}$ to fulfill $I_{A}\circ\tilde{\Phi}[I_{A},I_{B}]\circ\tilde{\Phi}[I_{A},I_{B}]\sim I_{B}$ instead of calculating the square root of a network’s output that directly registers $I_{A}$ to $I_{B}$. For each registration network, the output of $N_{\theta}[I_{A},I_{B}]-N_{\theta}[I_{B},I_{A}]$ was regularized using the selective Jacobian determinant regularization loss (Mok and Chung, 2020). The final output of the combined registration networks, was regularized using bending energy loss (Johnson and Christensen, 2001), with LNCC as the similarity metric. Additionally, we found that further smoothing the stationary velocity field (SVF) reduced the inverse consistency error, so we applied average pooling with a window size of 3 twice. During inference, the resulting displacement fields were upscaled to the original image resolution.
5.6 SSKJLBW: Unsupervised Single-Shot Registration
@@
To address the challenges of severe noise interference, large-scale deformations caused by respiratory motion, and limited training data in thoracic cone-beam CT (ThoraxCBCT) image registration for clinical radiotherapy, this paper proposes a single-shot unsupervised registration framework. Inheriting the optimization paradigm of GroupRegNet (Zhang et al., 2021), the method achieves robust registration without annotated data through multi-scale feature fusion and dynamic regularization constraints. The core workflow begins with noise-adaptive preprocessing of input images, followed by iterative optimization of deformation fields via a cascaded estimation module, and finally outputs smoothed deformation fields constrained by anatomical masks.
@@
The framework incorporates physics-guided preprocessing to decouple noise from anatomicalmdeformations. Input HU values are clamped to [0,2048] to suppress metal artifacts and extreme noise, while edge regions are zero-padded to mitigate respiratory motion artifacts. A lightweight 3D UNet processes standardized 128×128×64 image pairs, leveraging hierarchical features for deformation estimation—shallow layers enforce smooth large-scale deformations, while deeper layers refine local details.
@@ 6.1 ThoraxCBCT
The quantitative results for the ThoraxCBCT task are summarized in Table 3. While ConvexAdam achieved the highest overall ranking across all evaluation metrics, the official winner of the task was SynDeeds. Since ConvexAdam was provided by the organizers as a baseline method, it was not eligible for ranking. However, the performance of ConvexAdam and SynDeeds was very close across most metrics. An exemplary image pair is shown together with the warped moving image based on a displacement field achieved with ConvexAdam is shown in Fig. 5.
Figure 5: Exemplary image pair before registration (left: FBCT - moving image, middle: CBCT - fixed image) and after registration with ConvexAdam (right: FBCT - warped image)
When examining individual performance measures, SynDeeds achieved the highest mean Dice Score (DSC), while Fourier-Net obtained the lowest Target Registration Error (TRE). The best Hausdorff Distance 95th percentile (HD95) was observed for SSKJLBW, although this method exhibited a less smooth displacement field.
A more detailed breakdown of Dice coefficients achieved by all methods across different anatomical labels is shown in Fig. 6, while the distribution of Target Registration Errors is presented in Fig. 8.
@@ 6.2 OncoReg
For the initial challenge evaluation, we used a standard train-validation-test split. However, since many submitted methods did not rely on training data, we later expanded our evaluation to be more robust and meaningful. After completing additional annotations for all images, we conducted a three-fold cross-validation for all trainable models and evaluated the non-trainable methods on all additionally annotated images to ensure a fair and comprehensive comparison.
@@
The quantitative results are summarized in Table LABEL:tab:oncoreg and visualized in Fig. 8 and Fig. 9. More detailed results for bony structures (ribs and vertebrae) can be found in Appendix 11.
The official winner of the challenge was the method DINO-Reg Ensemble, which achieved the lowest target registration error and the highest Dice coefficient among all submissions. However, a late submission (TimH), which was not included in the initial ranking, outperformed DINO-Reg Ens. in terms of the 95th percentile of the Hausdorff Distance and the 30th percentile of the TRE, ultimately achieving the best overall performance.
These two leading methods were closely followed by Voxelmorph++ and ConvexAdam, while DeedsBCV and NiftyReg ranked lowest overall.
Figure 11 illustrates the relationship between TRE and DSC across all methods, while Figure 11 visualises the correlation between DSC and Hausdorff Distance for ten key anatomical structures. The lungs consistently exhibit high Dice scores, with the expected inverse relationship to HD95. In contrast, smaller structures such as the scapulae yield lower Dice scores. These patterns are also evident in Figure 9. Tumour registration reveals three distinct trends: high Dice scores are consistently associated with low HD95, but among the lower Dice scores, both low and very high HD95 values are observed. This behaviour differs from that seen in other anatomical structures.
@@
Initial 6.41 11.43 34.38 32.68
TimH 5.30 6.31 63.21 6.92 0.039 x 0.73
DINO-RegEns. 3.94 7.63 63.89 31.12 0.062 $>300$ 0.67
Voxelmorph++ 4.66 8.45 62.36 26.86 0.081 $<60$ 0.60
ConvexAdam 4.16 7.89 60.13 31.03 0.066 $(<5)$ 0.60
DINO-Reg 4.53 8.46 57.64 81.83 0.039 $>300$ 0.57
SynDeeds 4.51 7.93 59.93 28.29 0.100 x 0.52
MONAI SegResNet 4.78 8.97 53.63 32.37 0.048 x 0.46
Fourier-Net 6.01 10.52 60.67 23.53 0.103 $>300$ 0.44
DeedsBCV 7.99 12.28 58.33 22.33 0.154 $<15$ 0.34
Niftyreg 7.98 12.38 37.30 32.71 0.061 $<60$ 0.31
@@ 7.1 Evaluation Metrics and Annotation Challenges
Evaluating image registration remains a complex task, as standard metrics often fail to capture all aspects of registration quality and are highly dependent on annotation accuracy. In this challenge, the primary evaluation metric was the Target Registration Error, which relies on manually placed anatomical landmarks. However, landmark positioning is inherently uncertain in certain structures; while the carina served as a reliable reference point, the aortic arch and vertebral midpoints were more challenging to pinpoint precisely. To minimise inter-observer variability, all landmarks within a given image pair were annotated by the same medical expert, ensuring internal consistency. Significant effort was dedicated to refining these annotations, including extending the publicly available 4D-Lung dataset and comprehensively annotating our private dataset. This investment considerably enhanced the dataset’s usability. However, annotation challenges were particularly pronounced in cone-beam CT images, where lower resolution, poor contrast, and strong artifacts complicated precise landmark placement. In these cases, tools like TotalSegmentator proved invaluable, significantly reducing annotation time while maintaining quality in a semi-automated setting. Despite the comprehensive measures taken, the analysis of the registration results revealed that the label masks, particularly on the artefact-prone CBCT scans, often exhibited poorly smoothed surfaces. While the masks generally covered the approximate volume of the structures, they tended to be less reliable at the boundaries. This particularly affected the Dice coefficient and surface distance for small or thin structures, such as the clavicles and ribs. Furthermore, varying fields of view, as illustrated in Fig. 5, introduced an additional challenge for registration. As the CBCTs cover a more limited anatomical region compared to the FBCTs, which often include more extensive surrounding anatomy, many methods struggled to accurately register structures located near the image borders. This difficulty is reflected in the OncoReg task by the low Dice scores for the scapulae, lumbar and cervical vertebrae, as well as the most inferior and superior ribs, all of which are typically situated near the edges of the images. This trend is particularly evident in Fig. 12 and Fig. 13. Sample-based inspection further revealed that the notably high variability in Hausdorff distances was largely attributable to distortions at the image margins. The method TimH addresses this problem through extensive preprocessing: FBCT images are cropped to match the spatial extent of the corresponding CBCTs, ensuring alignment in the regions where both modalities contain anatomical information. This approach leads to significantly lower Hausdorff distances (see Table LABEL:tab:oncoreg). These insights are of high relevance for data preprocessing in future challenges.
@@ 7.2 Challenge Design and Implementation
Participants had the option to submit results via an automatic evaluation platform on Grand Challenge for the ThoraxCBCT task, but this feature was rarely utilised. Instead, most teams preferred local evaluation. In contrast, the provided container example was widely adopted, particularly for handling data loading and saving, which streamlined debugging in the Type 3 challenge setup. Despite these advantages, some recurring issues were identified that should be addressed in future Type 3 challenges.
Organizing such a challenge demands substantial time and resources, but efficient communication and accessibility between organisers and participants played a crucial role in ensuring a smooth workflow. Future iterations should build on this experience to optimise logistical aspects and further reduce participation barriers.
@@ 7.3 Key Findings and Methodological Insights
A key takeaway from the challenge is that the best-performing registration methods leveraged strong feature extraction strategies. The top-performing methods, ConvexAdam and DINO-Reg, demonstrate two distinct but effective approaches:
  • ConvexAdam combines handcrafted and semantic features within a two-stage optimisation process. The extracted semantic features play a crucial role in achieving robust registration across different modalities.
@@
Both methods prioritise feature extraction but differ in execution, ConvexAdam explicitly designs features, whereas DINO-Reg learns them in a self-supervised manner. Beyond these top approaches, an intriguing finding was the potential for further improving ConvexAdam’s performance by integrating it with VoxelMorph++. Specifically, using ConvexAdam-derived Förstner keypoints as input for VoxelMorph++ led to a refined registration process. This effect was particularly evident in OncoReg, where preprocessed keypoints significantly improved VoxelMorph++’s accuracy. However, in ThoraxCBCT, VoxelMorph++ performed notably worse than other methods. Further analysis revealed that its performance was hampered by suboptimal keypoint registration, which proved particularly detrimental in the weakly supervised setting. A post-challenge re-evaluation using keypoints registered by ConvexAdam showed substantial improvement—although this did not affect the official rankings. Another notable result was the enhancement of DeedsBCV by Li et al., leading to the superior performance of SynDeeds in both registration tasks. This demonstrates that traditional methods, when effectively adapted, can still compete with or even outperform newer deep-learning-based approaches.
@@ 7.4 Conclusion and Future Directions
[…]
The data for ThoraxCBCT is available on learn2reg.grand-challenge.org/oncoreg/. As described earlier the data of the OncoReg task can not be made publicly available, nevertheless, we are open for requests about evaluation on this data that will be computed in-house by us.
Appendix A Additional Results
Figure 10: OncoReg: Correlation between Dice Coefficient and Target Registration Error averaged over all structures and images per method.
Figure 11: OncoReg: Correlation between Dice Coefficient and 95th Percentile Hausdorff Distance for 10 main structures (soft tissue and bones, small and large organs).